40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Nov 13 01:03:28 2014, rana, jenne, Update, SUS, OL updates on ITMs and ETMs 
    Reply  Fri Nov 14 19:28:33 2014, Jenne, Update, SUS, ASS gain reduced for Yarm 
Message ID: 10708     Entry time: Thu Nov 13 01:03:28 2014     Reply to this: 10719
Author: rana, jenne 
Type: Update 
Category: SUS 
Subject: OL updates on ITMs and ETMs 

 We copied the new SRM filters over onto the OL banks for the ITMs and ETMs. We then adjusted the gain to be 3x lower than the gain at which it has a high frequency oscillation. This is the same recipe used for the SRM OL tuning.

Before this tune up, we also set the damping gains of the 4 arm cavity mirrors to give step response Q's of ~5 for all DOF and ~7-10 for SIDE.

ELOG V3.1.3-